OpenStudy (anonymous):

d+q=46 10d+25q=7.00 need help this is what i got so far and im stuck d+q=46 10d+25q=7.00 10(46-q)+25q=7.00 460-10q=25q=7.00 460+35q=7.00 I don't even know if it's right. help!!!

OpenStudy (anonymous):

yeah your equation is correct d+q = 46 which gives d = 46-q now use the dollar value given d*10 + q*25 = 700 so (46-q)*10 + q*25 = 700 Solve for q get d from q